# auto correlation matrix ADC analog to digital converter

### v1

@@ III-D KalmanCZT
III-E Performance Indicators
We use the Cramér-Rao bound (CRB) for benchmarking all sensing estimates. The CRB is a fundamental lower bound of the estimation error, but does not consider time dependence of the parameters to be estimated. In a tracking scenario, we may get an estimation accuracy below the CRB. For the AoA, the CRB can be formulated as
@@ IV-A Parameters
We consider an OFDM system with carrier frequency $f_{\text{c}}=5\,$GHz, a bandwidth of $B=25\,$MHz, and $N=2048$ subcarriers with a cyclic prefix length of $N_{\text{cp}}=30$. We group $M=259$ OFDM symbols in one frame for sensing. Therefore, the bounds (12)-(14) in the chosen scenario can be calculated as
@@
With the native RDM resolution, the range is the most limiting factor to reconstruct the trajectory accurately, as already indicated in Fig. 1.
@@ IV-B Configuration of Interpolation Methods
To get the same range resolution in the CZT as in the ZP, we configure the CZT in (7) at time step $k$ as
$A_{\text{v}}$ $=0,\quad W_{\text{v}}=\mathrm{e}^{\mathrm{j}2\pi M},\quad M=259\quad\text{and}$
$A_{\text{r}}$ $=\mathrm{e}^{\mathrm{j}2\pi\left(\frac{r_{k-1}}{Nr_{\text{res}}}-2^{-5}\right)},\quad W_{\text{r}}=\mathrm{e}^{\mathrm{j}\pi 2^{-14}},\quad N_{\text{CZT}}=2^{11}.$
The parameter ${r}_{k-1}$ denotes the range estimated in the previous time step.
@@ IV-C Configuration of Tracking Algorithms
with $\text{diag}(\cdot)$ denoting the diagonal matrix operator. We initialize ${\boldsymbol{\Sigma}}_{\text{est}}={\boldsymbol{0}}$ if the initial position is known and ${\boldsymbol{\Sigma}}_{\text{est}}=\text{diag}\begin{pmatrix}r_{\text{res}}&v_{\text{res}}&\frac{1}{2}\end{pmatrix}^{\top}$ if a measurement of the initial position is available. For the KalmanCZT, we choose the same initial ${\boldsymbol{\Sigma}}_{\text{meas}}$, ${\boldsymbol{\Sigma}}_{\text{est}}$, ${\boldsymbol{\Sigma}}_{\text{pred}}$ and prediction matrix ${\boldsymbol{F}}$.
@@ IV-E Performance Comparison
TABLE II: RMSE of resolution enhancement
@@
Assuming that the initial position is known, as commonly assumed in the tracking literature, the RMSE of the estimates for $1000$ trajectories consisting of $92$ points is shown in Tab. II. In addition to the estimates, we show the mean Euclidean distance between estimated and true positions in the last column. The range and velocity error and of the RDM and ZP match the bounds of (13), while the AoA error does not reach the CRB (12). This is expected behavior, since more intricate estimators such as superresolution methods are needed to achieve performance close to CRB.
@@
We achieve a very similar performance for all velocity estimators. We have configured the CZT and ZP to result in the same velocity resolution as the RDM. As the velocity can fall off quickly, tracking methods do not improve estimation and in the case of KalmanCZT even slightly worsen it, as the measurement tends to be more accurate than the prediction.
Additional tracking does not seem to improve the AoA estimate. Similarly to velocity tracking, the measurement is more accurate than the prediction.
@@
In Fig. 3, we show the RMSE of the range estimate for different estimation and tracking variants averaged over $1000$ trajectories, under the assumption that the initial ground truth values are known. Therefore, the error in the first time step is zero for all algorithms with tracking component. The estimate based on the native RDM performs worst and doesn’t change over time. The Kalman filter starts with an error of zero that gradually increases up to $\approx$0.250.250.25\,0.25m. The EBM starts with a behavior similar to the Kalman filter but reaches its stable error plateau after $15$ samples and maintains an error of 0.330.330.33\,0.33m for the rest of the observation. The performance of CZT and ZP is identical, except for the first time step where the CZT performance equals the RDM estimate, as it needs initialization of its observation window. The difference in the first time step is the reason for the higher RMSE of CZT in Tab. II when compared to ZP. Lastly, leveraging not only a tracking mean but also a tracking variance in the KalmanCZT, the performance jumps from the initial estimate with an error of 00 to an error plateau of 2.52.52.5\,2.5mm.
Results for the case where the initial position is not fully known, but instead the initial RDM estimate is used for all initial tracking estimates, are shown in Fig. 4. For the Kalman filter and the EBM, we observe a decreasing error behavior from the initial RDM resolution in contrast to Fig. 3. The EBM shows a consistent small gap to the performance of the Kalman filter. The KalmanCZT has the best performance, starting at at a slightly higher error but very quickly decreasing to the same error floor as in Fig. 3.
Figure 4: Tracking accuracy over time RMSE vs trajectory index with measured start position
@@ V Conclusion
In this paper, we provide an analysis on resolution boosting methods with a focus on range estimation in OFDM-based JCAS systems. We compare tracking methods and interpolation methods and propose the KalmanCZT approach, that combines estimation based on the CZT with Kalman filtering. The range estimation error can be reduced from $\approx$1.71.71.7\,1.7m to $\approx$333\,3mm and the importance of tracking initialization is pointed out by comparing tracking with a known initial position and a known initial estimate. We provide a detailed complexity analysis, noting that the additional complexity for tracking algorithms is very low compared to the computation of the RDM. Application of the CZT instead of the FFT yields improved resolution, for a controlled increase in computational complexity. The KalmanCZT approach provides a flexible resolution controlled by the dynamics of the tracked object, enabling efficient use of the computational resources.

### v4

@@ III-C KalmanCZT
III-D Performance Indicators
We use the Cramér-Rao bound (CRB) for benchmarking all sensing estimates. The CRB is a fundamental lower bound of the estimation error, but does not consider time dependence of the parameters to be estimated. In a tracking scenario, we may get an estimation accuracy below the CRB. For the AoA, the CRB can be formulated according to [13, Ch. 8.4] as
@@ IV-A Parameters
We consider an OFDM system with carrier frequency $f_{\text{c}}=5\,$GHz, a bandwidth of $B=25\,$MHz, and $N=2048$ subcarriers with a cyclic prefix length of $N_{\text{cp}}=30$. All simulation scenarios assume a constant $\mathrm{SNR}_{\varphi}=0\,$dB, independent of range variation. We group $M=259$ OFDM symbols in one frame for sensing. Therefore, the bounds (11)-(13) can be calculated as
@@
With the native RDM resolution, the range is the most limiting factor to reconstruct the trajectory accurately while the AoA can be estimated quite accurately, as illustrated in Fig. 1.
@@ IV-B Configuration of Interpolation Methods
To get the same range resolution in the CZT as in the ZP, we configure the CZT in (7) as
$A_{\text{v}}$ $=-1,\quad W_{\text{v}}=\mathrm{e}^{\mathrm{j}2\pi/M},\quad M=259\quad\text{and}$
$A_{\text{r}}$ $=\mathrm{e}^{\mathrm{j}2\pi\left(\frac{\hat{r}}{Nr_{\text{res}}}-2^{-5}\right)},\quad W_{\text{r}}=\mathrm{e}^{\mathrm{j}\pi 2^{-14}},\quad N_{\text{CZT}}=2^{11}.$
@@ IV-C Configuration of Tracking Algorithms
with $\text{diag}(\cdot)$ denoting the diagonal matrix operator. We initialize ${\boldsymbol{\Sigma}}_{\text{est}}={\boldsymbol{0}}$ if the initial position is known exactly and ${\boldsymbol{\Sigma}}_{\text{est}}=\text{diag}\begin{pmatrix}r_{\text{res}}&v_{\text{res}}&\frac{1}{2}\end{pmatrix}^{\top}$ when using the native RDM of the initial position is available. For the KalmanCZT, we choose the same initial ${\boldsymbol{\Sigma}}_{\text{meas}}$, ${\boldsymbol{\Sigma}}_{\text{est}}$, ${\boldsymbol{\Sigma}}_{\text{pred}}$ and prediction matrix ${\boldsymbol{F}}$.
@@ IV-E Performance Comparison
TABLE II: RMSE with known initial position
@@
Assuming that the initial position is known exactly, as commonly assumed in the tracking literature, the RMSE of the estimates for $1000$ trajectories consisting of $92$ points is shown in Tab. II. In addition to the estimates, we show the mean Euclidean distance between estimated and true positions in the last column. The range and velocity error and of the RDM and ZP match the bounds of (12), while the AoA error does not reach the CRB (11). This is expected behavior, since more intricate estimators such as super-resolution methods are needed to achieve performance close to CRB.
@@
We achieve a very similar performance for all velocity estimators. We have configured the CZT and ZP to result in the same velocity accuracy as the RDM. As the velocity can fall off quickly, tracking methods do not improve estimation and in the case of KalmanCZT even slightly worsen it. This is most probably due to mismatched ${\boldsymbol{\Sigma}}_{\text{meas}}$ and ${\boldsymbol{\Sigma}}_{\text{pred}}$, and can be remedied by further tuning. Additional tracking does not seem to improve the AoA estimate either.
@@
In Fig. 3, we show the RMSE of the range estimate for different estimation and tracking variants averaged over $1000$ trajectories, under the assumption that the initial ground truth values are known. Therefore, the error in the first time step is zero for all algorithms with tracking component. The estimate based on the native RDM performs worst and does not change over time. The Kalman filter starts with an error of zero that gradually increases up to $\approx$0.250.250.25\,0.25m. The EBM starts with a behavior similar to the Kalman filter but reaches its stable error plateau after $15$ samples and maintains an error of 0.330.330.33\,0.33m for the rest of the observation. The performance of CZT and ZP is identical, except for the first time step where the CZT performance equals the RDM estimate, as it needs initialization of its observation window. The difference in the first time step is the reason for the higher RMSE of CZT in Tab. II when compared to ZP. Lastly, leveraging not only a tracking mean but also a tracking variance in the KalmanCZT, the performance jumps from the initial estimate with an error of 00 to an error plateau of 2.52.52.5\,2.5mm.
Results for the case where the initial position is not fully known, but instead the initial RDM estimate is used for all initial tracking estimates, are shown in Fig. 4. For the Kalman filter and the EBM, we observe a decreasing error behavior from the initial RDM accuracy in contrast to Fig. 3. The EBM shows a consistent small gap to the performance of the Kalman filter. The KalmanCZT has the best performance, starting at at a slightly higher error but very quickly decreasing to the same error floor as in Fig. 3.
Figure 4: Tracking accuracy over time RMSE vs trajectory index with start position being the RDM estimate
@@ V Conclusion
In this paper, we provide an analysis on estimation accuracy enhancing methods with a focus on range estimation in OFDM-based JCAS systems. We compare tracking methods and interpolation methods and propose the KalmanCZT approach, that combines estimation based on the CZT with Kalman filtering. The range estimation error can be reduced from $\approx$1.71.71.7\,1.7m to $\approx$333\,3mm and the importance of tracking initialization is pointed out by comparing tracking with a known initial position and a known initial estimate. We provide a detailed complexity analysis, noting that the additional complexity for tracking algorithms is very low compared to the computation of the RDM. Application of the CZT instead of the FFT yields improved estimation, for a controlled increase in computational complexity. The KalmanCZT approach provides a flexible estimation controlled by the dynamics of the tracked object, enabling efficient use of the computational resources.
For future investigations we are planning to extend the approach in order to apply it to real measurement data.